a wire when bent into the form of a square encloses an area of 121 sq.cm. if the wire is bent in the form of a circle, then what would be the area?
Answer by nerdybill(7384)

a wire when bent into the form of a square encloses an area of 121 sq.cm. if the wire is bent in the form of a circle, then what would be the area?
Let x = length of one side square
then
x^2 = 121
x = 11 cm
.
Perimeter of square = 4*11 = 44 cm
circumference of circle = 44 cm
2(pi)r = 44
(pi)r = 22
r = 22/(pi)
.
area of circle:
(pi)r^2
(pi)(22/(pi))^2
(pi)(484/(pi)^2)
484/(pi)
154.14 square cm